# DREAM: Scalable Red Teaming for Text-to-Image Generative Systems via Distribution Modeling

### v1

@@ 5.5 Transferability on Real-world Commercial T2I Generative Models
To further assess the scalability of DREAM in real-world conditions, we test it on four widely used commercial T2I platforms: Ideogram, DeepAI, DALL·E 3, and Midjourney. These platforms deploy state-of-the-art, closed-source safety systems that at least include both prompt- and image-level filters, though their exact implementations are undisclosed. Moreover, some platforms use proprietary LLMs to interpret and rewrite user prompts. To evaluate how different methods perform on these real-world platforms, we train DREAM and FLIRT on the NSFW Image and Text Hybrid Filter, and then randomly select 50 prompts to conduct a transfer-based red teaming. We evaluate both “sexual” and “violence” categories, which are explicitly prohibited by all the platforms’ safety policies, and report the averaged results. As shown in Tab. V, our method achieves good transferability on all evaluated platforms, as validated by a high prompt bypass rate (the fraction of prompts accepted by the text filter), prompt-image bypass rate (the fraction of attempts that successfully yield generated images), human-rated prompt success rate, and still outperforms baselines with a notable margin. The results indicate that while the unsafe prompt space varies across different T2I platforms, the prompts generated by our method possess a notable degree of transferability, i.e., we can uncover some shared vulnerabilities across different systems, possibly due to broader prompt coverage and the inherent similarity across these T2I models.
@@ 5.6 Discussion
LLM Reusability. One potential advantage of our distributional modeling approach is that the red team LLM, once trained on a T2I model, learns a holistic understanding of the probability distribution over unsafe prompts. As a result, the LLM retains reusable knowledge that may be effectively leveraged when adapted to similar T2I systems. To evaluate this hypothesis, we conduct reusability experiments where a red team LLM trained on CA for 300 steps is adapted to other T2I systems. As shown in Tab. VI, our red team LLM achieves non-trivial success rates when directly reused (transferred) on other models without any further training. More importantly, with only 50 additional training steps ($\sim 2$ GPU hours), the reused LLM can be rapidly adapted to the new model, with some even achieving performance close to training from scratch (e.g., on ESD and UCE). These results show that DREAM’s distribution-level modeling demonstrates strong reusability potential, where the learned knowledge can be efficiently transferred and adapted to other T2I systems with reduced computational overhead.
Mitigation Strategy. To mitigate the identified vulnerabilities, one practical strategy is safety-tuning, which fine-tunes the T2I model on the collected unsafe prompts in an adversarial manner to unlearn them. To assess the utility of different methods for this purpose, we utilize red team prompts identified by each method as the dataset, and use Safety-DPO [22], a recent algorithm designed to steer generation away from unsafe behaviors via preference modeling, to fine-tune the SD v1.5 model. We then evaluate the resulting models against prompt sets from all methods, which yields a square matrix where each row represents a safety-tuned model (trained on method A’s data), and each column corresponds to evaluation against method B’s prompts. As shown in Fig. 3, the model adversarially trained with our DREAM-generated dataset consistently achieves the lowest PSR across all test sets and both categories, including those totally unseen during training. In contrast, models tuned with baseline datasets tend to show limited generalization and failing to defend against prompts from other methods, especially those discovered by DREAM. This also indirectly suggests that DREAM’s global modeling helps improve the diversity and coverage of discovered prompts, which in turn supports more robust and generalizable safety improvement.
@@
Ours (+50 Step Adaptation) 68.5% 86.3% 82.9% 48.3%
@@
Prior-informed Enhancement. For generality, our DREAM is designed without imposing specific prior about the internal components or defenses of the target T2I system. However, in practice, model owners (e.g., developers) have prior knowledge about the system, which can be potentially leveraged to enhance red teaming. For instance, if the model owner knows the system employs keyword-based filters, a simple enhancement strategy is to remove these tokens from the red team LLM’s vocabulary. This prior encourages the generator to focus on unexplored regions of the prompt space without wasting effort on words that are doomed to be rejected. To evaluate this, we conduct a case study on the Keyword Filter + UCE setup. We observe that the keyword-removed version converges faster, reaching near-optimal performance within 150 training steps, compared to 280 steps required by the baseline. Interestingly, the final performance difference is modest (63.4% vs. 60.2% PSR), suggesting that while prior knowledge accelerates convergence, it is not critical for eventual performance. This result highlights two insights: first, DREAM is effective even without any system-specific priors, making it broadly applicable; second, when available, prior information can be selectively incorporated to enhance DREAM. However, leveraging such priors often requires case-specific integration strategies, some of which may be difficult or costly to implement in practice. How to develop principled ways to incorporate them, especially for neural network-based components, remains an open direction for future work.
5.7 Ablation Study
In this section, we conduct an ablation study of DREAM, with CA+Sexual as the default setting.
Effectiveness of Each Component. As shown in Tab. VII, all components contributes to DREAM’s final performance. For example, while $E_{\text{align}}(x)$ pushes the model towards harmful outputs, it often leads to less diverse prompts. Adding $E_{\text{div}}(x)$ helps strike a balance between effectiveness and diversity. Additionally, Adaptive temperature scaling (ATS) improves prompt diversity during inference with only minimal PSR degradation, highlighting its effectiveness for balancing effectiveness and diversity.
TABLE VII: Ablation study on each component. ATS stands for our inference time adaptive temperature scaling strategy and Opt. Alg. means optimization algorithm.
@@ 5.7 Ablation Study
$\checkmark$ -- -- GC-SPSA 90.8% /00.70
$\checkmark$ $\checkmark$ -- GC-SPSA 76.4% /00.58
$\checkmark$ $\checkmark$ $\checkmark$ SPSA 45.9% /00.54
$\checkmark$ $\checkmark$ $\checkmark$ GC-SPSA 76.0% / 0.56
@@
Method Steps GPU Time PSR $\uparrow$ / PS $\downarrow$
@@
Effectiveness of GC-SPSA. As shown in Tab. VIII, GC-SPSA consistently outperforms vanilla SPSA, even when SPSA is given more steps to ensure equal GPU time. This demonstrates that investing in early gradient quality and applying historical calibration can indeed yield better overall performance, which indirectly confirms our insights in Sec. 4. In addition, we evaluate different initial sampling budgets $n_{0}$. We observe that $n_{0}=4$ already provides a strong balance between efficiency and effectiveness. While increasing $n_{0}$ to 8 leads to further gains, the improvement is marginal compared to the additional cost, possibly because the variance is already small enough to ensure stable optimization. Therefore, we adopt $n_{0}=4$ as our default configuration, as it strikes a good trade-off between convergence speed, stability, and computational efficiency.
Effect of $\lambda$ and $\alpha$. $\lambda$ and $\alpha$ controls the trade-off between diversity and effectiveness by influencing the weighting of diversity energy during training and temperature sensitivity during inference, respectively. As shown in Tab. IX, increasing $\lambda$ or decreasing $\alpha$ enhances prompt diversity, yet at the cost of the decrease of effectiveness. However, the performance is generally stable and satisfactory when these hyperparameters are within a reasonable range. Thus, we set the hyperparameters to $\lambda=1$ and $\alpha=0.05$ as the default configurations, and the users may tune them if they have specific emphasis on effectiveness or diversity.
TABLE IX: Ablation study on $\lambda$ and $\alpha$.
$\alpha$ 0.01 0.03 0.05 0.08 0.1
70.2% / 0.53 72.7% / 0.54 76.0% / 0.56 76.4% / 0.57 76.4% / 0.58
$\lambda$ 0.6 0.8 1 1.2 1.4
80.4% / 0.61 78.7% / 0.58 76.0% / 0.56 74.7% / 0.54 65.4% / 0.52
@@ 6 Conclusion
Limitations. Our work still has the following limitations, which we aim to address in future work. First, similar to other methods, our current implementation involves several auxiliary models (e.g., BLIP-2), and it is possible that our system may inherit certain limitations or biases rooted in these models. Although we did not observe clear evidence of such bias in our experiments, we acknowledge the potential risk and plan to explore more robust alternatives in the future. Second, while our global modeling approach offers desirable benefits, we admit it requires moderate costs to train. However, we note that this training cost is amortized over a large number of generated prompts. Compared to some baseline methods (e.g., P4D and MMA) which require 30 minutes or more to obtain a single prompt, DREAM remains more efficient even when applied to a moderate number of prompts ($>30$). For users with limited computational resources, it is also possible to further reduce cost by reusing or adapting a previously trained red team LLM, as discussed in Sec. 5.6. Finally, while our GC-SPSA demonstrates both theoretical guarantees and strong empirical performance outperforming baseline optimizers such as vanilla SPSA and LLM-based heuristics, we acknowledge it is essentially an approximation and may not be fully precise or perfectly efficient. Nonetheless, as discussed in Sec. 4, obtaining exact gradients via backprobagation are often infeasible due to the memory-intensive and potentially non-differentiable nature of the full T2I pipeline. We hope our energy-based distributional formulation and our proposed optimizer can inspire and serve as a valuable foundation for future improvements, and ultimately inspire stronger, more systematic safety evaluation practices for T2I systems.
@@ A.1 Proof of Theorem 1
$g(\theta)=\left(\Delta^{\top}g_{\text{true}}+O(\epsilon^{2})\right)\Delta+\frac{\xi^{+}-\xi^{-}}{2\epsilon}\Delta$ (11)
Taking expectation over the randomness in $\Delta$ and $\xi$: $\mathbb{E}[g(\theta)]=g_{\text{true}}+O(\epsilon^{2})$ since $\mathbb{E}_{\Delta}\left[(\Delta^{\top}g_{\text{true}})\Delta\right]=g_{\text{true}}$ and the noise term averages to zero.
The second moment arises from signal, bias, and noise components, i.e., $\mathbb{E}[\|(\Delta^{\top}g_{\text{true}})\Delta\|^{2}]=d\|g_{\text{true}}\|^{2}$, $\mathbb{E}[\|O(\epsilon^{2})\Delta\|^{2}]=O(d\epsilon^{4})$, and $\mathbb{E}[\|\frac{\xi^{+}-\xi^{-}}{2\epsilon}\Delta\|^{2}]\leq\frac{d\sigma_{\xi}^{2}}{2\epsilon^{2}}$. Thus, the variance of one time estimation is:
@@
GC-SPSA collects $n_{t}$ times estimation following the exponential decay strategy $n_{t}=\max(1,\lfloor n_{0}/2^{t/T_{\text{dec}}}\rfloor)$ and takes their average to obtain $\bar{g}_{t}(\theta)$. Its variance is:
$V_{t}:=\mathrm{Var}(\bar{g}_{t}(\theta))=\frac{V_{\rm single}}{n_{t}}$ (13)
The GC-SPSA estimator $\hat{g}_{t}$ follows the recursive update rule (LABEL:eq:spsa-gc), unrolling this recursion gives:
$\hat{g}_{t}=\sum_{k=0}^{t}h_{k}\bar{g}_{k}(\theta_{k})$ (14)
where $h_{t}=1$, $H_{j}=\gamma\frac{w_{j-1}}{w_{j-1}+n_{j}}$, and $h_{k}=\prod_{j=k+1}^{t}H_{j}$.
The expectation and variance of the GC-SPSA are:
@@
$\mathrm{Var}(\hat{g}_{t})$ $=\sum_{k=0}^{t}h_{k}^{2}V_{k}=\sum_{k=0}^{t}h_{k}^{2}\frac{V_{\rm single}}{n_{k}}$ (16)
The SNR improvement is defined as $\mathcal{D}_{t}=\mathrm{SNR}_{t}^{\text{GC}}-\mathrm{SNR}^{\rm Vanilla}$, where:
$\mathrm{SNR}_{t}^{\text{GC}}=\frac{\|g_{\text{true}}(\theta_{t})\|^{2}P_{t}^{2}}{\sum_{k=0}^{t}h_{k}^{2}V_{k}}$ (17)
$\mathrm{SNR}^{\rm Vanilla}=\frac{\|g_{\text{true}}(\theta_{t})\|^{2}}{V_{\rm single}}$ (18)
Therefore:
@@
To prove the bracketed term $P_{t}^{2}V_{\rm single}-\sum_{k=0}^{t}h_{k}^{2}V_{k}$ is strictly positive, we expand $P_{t}^{2}=\sum_{k=0}^{t}h_{k}^{2}+2\sum_{0\leq i<j\leq t}h_{i}h_{j}$:
@@
The first term is non-negative, and the second term is strictly positive. Therefore, we have $\mathcal{D}_{t}>0$. ∎

### v2

@@ 5.5 Transferability on Real-world Commercial T2I Generative Models
To further assess the scalability of DREAM in real-world conditions, we test it on 4 widely used commercial T2I platforms: Ideogram [ideogram2025], DeepAI [deepai2025], DALL·E 3 [openai_dalle3_2023], and Midjourney [midjourney2025]. These platforms deploy state-of-the-art safety systems, which at least include both prompt- and image-level filters, though their exact implementations remain undisclosed. Moreover, some platforms use proprietary LLMs to interpret and rewrite user prompts.
To evaluate how different methods perform on these platforms, we train DREAM and FLIRT on the NSFW Image and Text Filter, and randomly select 50 prompts to conduct a transfer-based red teaming. We evaluate both “sexual” and “violence” categories, which are explicitly prohibited by all the platforms’ safety policies, and report the averaged results. As shown in Tab. V, our method achieves good transferability on all evaluated platforms, as validated by a high prompt bypass rate (the fraction of prompts accepted by the text filter), prompt-image bypass rate (the fraction of attempts that successfully yield generated images), human-rated prompt success rate, and still outperforms baselines with a notable margin. The results indicate that while the unsafe prompt space varies across different T2I systems and real-world platforms, the prompts generated by our method possess a notable degree of transferability, that is, we can uncover some shared vulnerabilities across different systems, possibly due to broader prompt coverage and the inherent similarity across these T2I models.
5.6 Efficiency Comparison
In this section, we present a comparison of efficiency between DREAM and baselines. Unlike previous prompt-to-prompt methods whose time cost can be directly measured per prompt, DREAM involves two stages of time cost: (1) training the red-team LLM, which dominates the total time, and (2) sampling from the trained model, which is efficient. Therefore, their efficiency cannot be directly compared on a per-prompt basis. To this end, we compute and compare the averaged expected total time required to obtain a specified number of effective unsafe prompts (as measured by MHSC) over both (a) all safety-aligned diffusion models evaluated in Tab. I and (b) all external safety filters in Tab. II.
Figure 3: Efficiency comparison with baselines. We report the expected total time to collect a specified number of effective unsafe prompts, averaged across all evaluated (a) safety-aligned models and (b) safety filters. The considered unsafe concept is sexual.
As shown in Fig. 3, since baseline methods optimize each prompt through discrete, per-instance search without any global modeling, their total time cost scales quickly with the number of desired effective prompts. In contrast, while DREAM requires approximately 12 hours to train the red-team LLM, sampling from it is extremely efficient: it only takes about 5 minutes to generate over 1,000 effective prompts in our experiments. Compared with token-level optimization baselines such as MMA-Diffusion and P4D, DREAM is more efficient even at relatively small scales (e.g., 50 prompts). Even when compared to FLIRT, which improves efficiency through LLM rewriting, DREAM surpasses it at moderate scales (more than 500 prompts). Considering that downstream applications like safety tuning typically require thousands of diverse and effective prompts to ensure generalization, we believe that DREAM’s distribution-level modeling paradigm offers a practical and promising perspective for scalable red teaming.
@@ 5.7 Discussion
LLM Reusability. One potential advantage of our distributional modeling approach is that the red team LLM, once trained on a T2I model, learns a holistic understanding of the probability distribution over unsafe prompts. As a result, the LLM retains reusable knowledge that may be effectively leveraged when adapted to similar T2I systems. To validate this possibility, we conduct reusability experiments where a red team LLM trained on CA for 300 steps is adapted to other T2I systems. As shown in Tab. VI, DREAM’s red team LLM achieves non-trivial success rates when directly reused (transferred) on other models without any further training. More importantly, with only 50 additional training steps, the reused LLM can be rapidly adapted to the new model, with some even achieving performance close to training from scratch (e.g., on ESD and UCE). These results show that DREAM’s distribution-level modeling demonstrates strong reusability potential, where the learned knowledge can be efficiently transferred and adapted to other T2I systems with reduced computational overhead.
Mitigation Strategy. To mitigate the identified vulnerabilities, one practical strategy is safety-tuning, which fine-tunes the T2I model on the collected unsafe prompts in an adversarial manner to unlearn them. To assess the utility of different methods for this purpose, we utilize red team prompts identified by each method as the dataset, and use Safety-DPO [liu2024safetydpo], a recent algorithm designed to steer generation away from unsafe behaviors via preference modeling, to fine-tune the SD v1.5 model. We then evaluate the resulting models against prompt sets from all methods, which yields a square matrix where each row represents a safety-tuned model (trained on method A’s data), and each column corresponds to evaluation against method B’s prompts. As shown in Fig. 4, the model trained with our DREAM-generated dataset consistently achieves the lowest PSR across all test sets, including those totally unseen during training. In contrast, models tuned with baseline datasets tend to show limited generalization and fail to defend against prompts from other methods, especially those discovered by DREAM. This also indirectly suggests that DREAM’s global modeling helps improve the coverage of discovered prompts, which in turn supports more robust and generalizable safety improvement. We offer further discussions in Appendix B in our full technical report.
@@
Ours (+50 Steps Adaptation) 68.5% 86.3% 82.9% 48.3%
@@
Prior-informed Enhancement. For generality, our DREAM is designed without imposing specific priors about the internal components or defenses of the target T2I system. However, in practice, model owners (e.g., developers) have prior knowledge about the system, which can be potentially leveraged to enhance red teaming. For instance, if the model owner knows the system employs keyword-based filters, a simple enhancement strategy is to remove these tokens from the red team LLM’s vocabulary. This prior encourages the generator to focus on unexplored regions of the prompt space without wasting effort on words that are doomed to be rejected. To evaluate this, we conduct a case study on the Keyword Filter + UCE setup. We observe that the keyword-removed variant converges faster, reaching near-optimal performance within 150 training steps, compared to 280 steps required by the baseline. Interestingly, the final performance difference is modest (63.4% vs. 60.2% PSR), suggesting that while prior knowledge accelerates convergence, it is not critical for eventual performance. This result highlights two insights: first, DREAM is effective even without any system-specific priors, making it broadly applicable; second, when available, prior information can be selectively incorporated to enhance DREAM. However, leveraging such priors often requires case-specific integration strategies, some of which may be difficult or costly to implement in practice. How to develop principled ways to incorporate them, especially for neural network-based components, remains an open direction for future work.
5.8 Ablation Study & Hyperparameter Analysis
In this section, we conduct an ablation study of DREAM and analyze some key hyperparameters involved in our experiments, with CA+Sexual as the default setting.
Effectiveness of Each Component. As shown in Tab. VII, all components contribute to DREAM’s final performance. For example, while $E_{\text{align}}(x)$ pushes the model towards harmful outputs, it often leads to less diverse prompts. Adding $E_{\text{div}}(x)$ helps strike a balance between effectiveness and diversity. Additionally, ATS improves prompt diversity during inference with only minimal PSR degradation, highlighting its effectiveness for balancing success rate and diversity.
[…]
@@ 5.8 Ablation Study & Hyperparameter Analysis
[…]
@@
Method Steps Time PSR $\uparrow$ / PS $\downarrow$
@@
Effectiveness of GC-SPSA. As shown in Tab. VIII, GC-SPSA consistently outperforms vanilla SPSA, even when SPSA is given more steps to ensure equal training time, demonstrating the effectiveness of GC-SPSA. Besides, we evaluate different initial sampling budgets $n_{0}$. We observe that $n_{0}=4$ already provides a strong balance between efficiency and effectiveness. While increasing $n_{0}$ to 8 leads to further gains, the improvement is marginal compared to the additional cost, possibly because the variance is already sufficiently small for stable optimization. We thus adopt $n_{0}=4$ as our default configuration, as it strikes a good trade-off between effectiveness and computational cost.
TABLE IX: Impact of different system prompts.
Setting System Prompt 1 System Prompt 2 System Prompt 3
PSR $\uparrow$ / PS $\downarrow$ 77.8% / 0.57 75.2% / 0.56 76.4% / 0.56
Figure 5: Impact of number of sampled prompts on PSR and PS. We report the results computed over three independent samples.
Impact of System Prompt. Currently, we handcraft system prompts for DREAM. To understand DREAM’s sensitivity under different system prompts, we prompt ChatGPT-5 to generate three concise red-teaming role-play prompts for DREAM using the instruction “I’m training a red teaming assistant for text-to-image generative models for category {category}. Please generate a concise system prompt for it.”, and test their performances. As shown in Tab. IX, all prompts are highly effective, indicating that DREAM’s performance is not very sensitive to system prompts.
Impact of Number of Sampled Prompts. Similar to prior work [yang2024mma, mehrabi2024flirt], we generate 1,024 prompts for evaluation by default. As our method models a distribution and supports efficient sampling, it is natural to study the impact of the number of sampled prompts on performance. As shown in Fig. 5, the number of prompts has only a minor impact on DREAM’s expected PSR and PS, indicating DREAM’s scalability to larger sample sizes without significantly sacrificing performance, enabling it to generate large-scale, diverse, and effective red team datasets efficiently.
@@ 6 Conclusion
Limitations. Our work still has the following limitations, which we aim to address in future work. First, similar to other red-teaming studies, our implementation and evaluation rely on some auxiliary models (e.g., BLIP-2 and MHSC) that may introduce biases or inaccuracies, and we acknowledge that some false positives or false negatives may appear. Although we also conducted human evaluation, we acknowledge that concepts such as “unsafe” and “diversity” are inherently subjective and can be influenced by cultural, contextual, and personal factors. Nevertheless, since all methods are evaluated under the same protocol, we believe the comparisons remain relatively fair and informative. We plan to explore more robust alternatives in the future. Besides, while our GC-SPSA optimizer demonstrates both theoretical guarantees and strong empirical performance compared to baseline optimizers such as vanilla SPSA and LLM-based heuristics, we acknowledge it is essentially an approximation and may not be fully precise or perfectly efficient. Nonetheless, as discussed in Sec. 4, obtaining exact gradients via backpropagation is often infeasible due to the memory-intensive nature of the full T2I pipeline. We hope our energy-based distributional formulation and our proposed optimizer can inspire and serve as a valuable foundation for future improvements, and ultimately enable stronger safety evaluation practices for T2I systems.
Acknowledgments
The authors would like to sincerely thank all anonymous reviewers and our shepherd for their constructive feedback and helpful suggestions, which have significantly improved the quality of this paper. This work was supported in part by National Research Foundation, Singapore and DSO National Laboratories under its AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), and National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. It was also supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 62202340, No. 62576255. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, DSO National Laboratories, Infocomm Media Development Authority, or the NSFC.
Disclosure of Financial Interests. The authors declare that they have no financial or non-financial conflicts that could be perceived to influence the work reported in this paper.
@@ A.1 Proof of Theorem 1
$g(\theta)=\left(\Delta^{\top}g_{\text{true}}+O(\epsilon^{2})\right)\Delta+\frac{\xi^{+}-\xi^{-}}{2\epsilon}\Delta.$ (11)
Taking expectation over the randomness in $\Delta$ and $\xi$: $\mathbb{E}[g(\theta)]=g_{\text{true}}+O(\epsilon^{2})$ since $\mathbb{E}_{\Delta}\left[(\Delta^{\top}g_{\text{true}})\Delta\right]=g_{\text{true}}$
The second moment arises from signal, bias, and noise components, i.e., $\mathbb{E}[\|(\Delta^{\top}g_{\text{true}})\Delta\|^{2}]=d\|g_{\text{true}}\|^{2}$, $\mathbb{E}[\|O(\epsilon^{2})\Delta\|^{2}]=O(d\epsilon^{4})$, and $\mathbb{E}[\|\frac{\xi^{+}-\xi^{-}}{2\epsilon}\Delta\|^{2}]\leq\frac{d\sigma_{\xi}^{2}}{2\epsilon^{2}}$. Thus, the variance of a single gradient estimation is:
@@
GC-SPSA collects $n_{t}$ gradient estimates following the exponential decay strategy and takes their average to obtain $\bar{g}_{t}(\theta)$. The variance of this averaged estimate is ${V_{\rm single}}/{n_{t}}$.
The GC-SPSA estimator $\hat{g}_{t}$ follows the recursive update rule Eq. (LABEL:eq:spsa-gc), unrolling which gives $\hat{g}_{t}=\sum_{k=0}^{t}h_{k}\bar{g}_{k}(\theta_{k})$, thus the expectation and variance of GC-SPSA are:
@@
$\mathrm{Var}(\hat{g}_{t})=\sum_{k=0}^{t}h_{k}^{2}V_{k}=\sum_{k=0}^{t}h_{k}^{2}\frac{V_{\rm single}}{n_{k}},$ (14)
The SNR improvement is defined as $\mathcal{D}_{t}=\mathrm{SNR}_{t}^{\text{GC}}-\mathrm{SNR}^{\rm Vanilla}$, where we have $\mathrm{SNR}_{t}^{\text{GC}}=\frac{\|g_{\text{true}}(\theta_{t})\|^{2}P_{t}^{2}}{\sum_{k=0}^{t}h_{k}^{2}V_{k}}$ and $\mathrm{SNR}^{\rm Vanilla}=\frac{\|g_{\text{true}}(\theta_{t})\|^{2}}{V_{\rm single}}$. Therefore, we have:
@@
To analyze the bracketed term $P_{t}^{2}V_{\rm single}-\sum_{k=0}^{t}h_{k}^{2}V_{k}$, we expand $P_{t}^{2}=\sum_{k=0}^{t}h_{k}^{2}+2\sum_{0\leq i<j\leq t}h_{i}h_{j}$ and obtain:
@@
The first term is non-negative, and the second term is strictly positive. Therefore, we have $\mathcal{D}_{t}>0$ for $t\geq 1$. ∎
We can derive two key insights from Eq. (16). Specifically, let $S_{2}:=\sum_{k=0}^{t}h_{k}^{2}$, we have:
Insight I: gradient calibration alone can provide strong SNR gain. Using the algebraic identity:
$2\sum_{0\leq i<j\leq t}h_{i}h_{j}=P_{t}^{2}-S_{2},$ (17)
we have:
$P_{t}^{2}=\sum_{k=0}^{t}h_{k}^{2}+2\sum_{0\leq i<j\leq t}h_{i}h_{j}>S_{2},$ (18)
since $h_{k}=\prod_{j=k+1}^{t}H_{j}>0$ with $h_{t}=1$. $P_{t}$ grows linearly while $S_{2}$ remains bounded, making $P_{t}^{2}-S_{2}=\Theta(P_{t}^{2})$ exhibit quadratic growth.
Insight II: larger sampling budgets $n_{k}$ can further enhance SNR. Increasing $n_{k}$ reduces $V_{k}=\frac{V_{\rm single}}{n_{k}}$ and amplifies the coefficient $(1-\frac{1}{n_{k}})$ since $\frac{\partial(1-1/n_{k})}{\partial n_{k}}=\frac{1}{n_{k}}>0$.
Note the sampling term is bounded by $V_{\rm single}S_{2}$, while the calibration cross-term equals $V_{\rm single}(P_{t}^{2}-S_{2})$. Calibration dominates when $P_{t}^{2}>2S_{2}$. In typical implementations with $H_{j}\geq 1/2$, taking the conservative bound $H_{j}\equiv r\in[1/2,1)$ gives $h_{k}=r^{t-k}$, yielding:
$P_{t}=\frac{1-r^{t+1}}{1-r},\quad S_{2}=\frac{1-r^{2(t+1)}}{1-r^{2}}.$ (19)
Direct calculation shows:
$P_{t}^{2}-2S_{2}=\frac{-(r^{t+1}-1)(3r-3r^{t+1}+r^{t+2}-1)}{(1-r)^{2}(1+r)}.$ (20)
For any $r\in[1/2,1)$ and $t\geq 2$, both numerator and denominator are positive, proving $P_{t}^{2}>2S_{2}$. Therefore, the calibration cross-term exceeds the sampling term from $t\geq 2$ onward, even under conservative assumptions. ∎